the radiator in a certain make care needs to contain 40 liters of 40% antifreeze. the radiator now contains 40 liters of 20% antifreeze. how many liters of this solution must be drained and replaced with 100% antifreeze to get the desired strength?
Answer by stanbon(75887)

the radiator in a certain make care needs to contain 40 liters of 40% antifreeze. the radiator now contains 40 liters of 20% antifreeze. how many liters of this solution must be drained and replaced with 100% antifreeze to get the desired strength?
----------------------------
Equation:
alcohol - alcohol + alcohol = alcohol
0.20*40 - 0.20x + 1.00x = 0.40*40
---
Multiply thru by 100 to get:
20*40 - 20x + 100x = 40*40
80x = 20*40
x = 10 liters (amt. of liquid to drain and replace_
==========
Cheers,
Stan H.
==========